# Peter Reinhard Hansen§    and   Chen Tong

### v3

@@ 2 A New Method for Computing Moments (CMGF)
Let $X$ satisfy Assumption 1 and let $\xi\in\mathbb{R}$. For $k\in\mathbb{N}$ we have
$\mathbb{E}[(X-\xi)^{k}]$ $=$ $\frac{k!}{\pi}\int_{0}^{\infty}\operatorname{Re}\left[\frac{e^{-\xi z}M_{X}(z)+(-1)^{k}e^{\xi z}M_{X}(-z)}{z^{k+1}}\right]\mathrm{d}t,$ (6)
where $z=s+it$ with $s\in\mathcal{N}_{+}$.
Because $k$ is real, we have expressed the integral on the form, $\frac{k!}{\pi}\int_{0}^{\infty}\operatorname{Re}[\cdot]\mathrm{d}t$, which tends to be simpler to evaluate numerically than the equivalent expression using the form, $\frac{k!}{2\pi}\int_{-\infty}^{+\infty}[\cdot]\mathrm{d}t$.
@@
In some situations, Theorems 1, 2, and 3, provide different expressions for the same moments. For instance, (4) and (1) must agree for a non-negative random variable, because $\mathbb{E}|X|^{r}=\mathbb{E}[X^{r}]$ in this case, and for integer moments, these expressions must also agree with (6). This is indeed the case, because the additional term in (1) and the additional term in (6) are both zero for non-negative variables. This is an implication of Lemma A.2, which establishes that $\int_{-\infty}^{\infty}\frac{e^{-\left(s+it\right)x}}{\left(s+it\right)^{r+1}}\mathrm{d}t=0$ for all $x>0$ and all $s>0$.
We can combine Theorems 1 and 3 to compute odd moments over tail events.
Corollary 1 (Tail Moments).
Let $X$ satisfy Assumption 1, $\xi\in\mathbb{R}$, and let $k$ be odd number. Then
$\mathbb{E}[(X-\xi)^{k}1_{\{X>\xi\}}]$ $=$ $+\frac{k!}{\pi}\int_{0}^{\infty}{\rm Re}\left[\frac{e^{-\xi z}M_{X}(z)}{z^{k+1}}\right]\mathrm{d}t,$
$\mathbb{E}[(X-\xi)^{k}1_{\{X<\xi\}}]$ $=$ $-\frac{k!}{\pi}\int_{0}^{\infty}{\rm Re}\left[\frac{e^{-\xi z}M_{X}(z)}{z^{k+1}}\right]\mathrm{d}t,$
where $z=s+it$ for any $s\in\mathcal{N}_{+}$.
We can also compute the cdf, quantiles, and expected shortfall using the new moment expressions. We can obtain the cdf for a continuous random variable, $X$, from the MGF by Gil-Pelaez inversion,
$F_{X}(x)=\frac{1}{2}-\frac{1}{\pi}\int_{0}^{\infty}{\rm Re}\left[\frac{M_{X}(it)e^{-itx}}{it}\right]\mathrm{d}t,$
where we used that $M_{X}(it)$ is the characteristic function for $X$. Now the $\alpha$-quantile is given by $\xi_{\alpha}\equiv F_{X}^{-1}(\alpha)$, and we can use Corollary 1 (with $k=1$) to compute,
$\operatorname{ES}_{\alpha}(X)=-\frac{1}{\alpha}\mathbb{E}[X1_{\{X<\xi_{\alpha}\}}]=-\frac{1}{\alpha}\mathbb{E}[(X-\xi_{\alpha})1_{\{X<\xi_{\alpha}\}}]-\xi_{\alpha},$
which is known as the expected shortfall.
@@ 2.1 Example: Moments of NIG Distribution
To illustrate the new method, we use Theorem 1 to compute fractional moments of the normal-inverse Gaussian (NIG) distribution. The CMGF method is significantly faster and more accurate than both simulation-based methods and direct integration using the NIG density.
@@
The NIG distribution was introduced in Barndorff-Nielsen, (1978) and has four parameters, $\lambda$ (location), $\delta$ (scale), $\alpha$ (tail heaviness), and $\beta$ (asymmetry). It has density
@@
$f(x)=\tfrac{\alpha\delta}{\pi\sqrt{\delta^{2}+(x-\lambda)^{2}}}K_{1}\left(\alpha\sqrt{\delta^{2}+(x-\lambda)^{2}}\right)e^{\delta\gamma+\beta(x-\lambda)},\qquad x\in\mathbb{R}$
@@
where $\gamma=\sqrt{\alpha^{2}-\beta^{2}}$ and $K_{1}(\cdot)$ is the modified Bessel function of the second kind, and the MGF is given by $M_{X}(z)=\exp\left(\lambda z+\delta\left[\gamma-\sqrt{\alpha^{2}-(\beta-z)^{2}}\right]\right),$ for all $z\in\mathbb{C}$. We consider two standardized NIG distribution (zero mean and unit variance), which can be characterized by the two parameters, $(\xi,\chi)$, for $10\leq|\chi|<\xi<$ 1, where $\lambda=-\tfrac{\chi}{\xi}\zeta$, $\delta=\xi^{2}\sqrt{\xi^{2}-\chi^{2}}\zeta$, $\alpha=\xi\zeta$, and $\beta=\chi\zeta$, with $\zeta=\sqrt{1-\xi^{2}}/(\xi^{2}-\chi^{2})$, see Barndorff-Nielsen et al., (1985). Specifically we consider $(\xi,\chi)=(1/2,-1/3)$ and $(\xi,\chi)=(1/8,-1/16)$, the corresponding densities are shown in the left panel of Figure 1. The corresponding MGFs, $t\mapsto\operatorname{MGF}(s+it)$ are shown in the right panel for $s=1$ – solid lines for the real part and dashed lines for the complex part.
[…]
@@
We compute the absolute moments for $-0.85\leq r\leq 4.2$ for both distributions using Theorem 1 and the moments are shown in the upper panel of Figure 2. The solid black dot is the second (absolute) moment, which is known to be one for these standardized distributions, and the solid blue and red dots represent the fourth (absolute) moment, which are known to equal $3(1+4\chi^{2})/(1-\xi^{2})$. The values provided by the CMGF method are very accurate. The first 12 digits of the fourth moments are correct using a simple implementation in Julia.
@@
Figure 2: Absolute moments of the standardized NIG distribution with $(\xi,\chi)=(1/2,-1/3)$ (blue lines) and $(\xi,\chi)=(1/8,-1/16)$ (red lines). These are shown in the left panels, where dots are represent the known moments for $r=2$ and $r=4$. The middle panel is a snippet of the left panel which can be compared to simulation-based moments for 15 values of $r$. For each $r$ we computed $\mathbb{E}|X|^{r}$ 100 times as represented by the x-crosses. Each of these simulation-based moments are based on $N=$1,000,000 draws of NIG distribution. The right panel shows how many decimal places are accurate with a one standard deviation simulation error.
@@
Additionally, we compute the absolute moments using $N=1,000,000$ independent draws from the two NIG distributions. The simulation based moment is given by $\hat{\mu}_{N}^{r}=\frac{1}{N}\sum_{j=1}^{N}|X_{i}|^{r}$ and we estimate the $\sigma^{2}(r)=\operatorname{var}(|X_{j}|^{r})$ with 100 million draws. The accuracy of a simulated moment can be computed from its standard error. Assuming an error equal to one standard deviation, then the number of accurate decimal places is given by $-\log_{10}(\sigma(r)/\sqrt{N})$. [^4] For instance, an error where the first three decimal places are zero, say $0.00099$, will map to $-\log_{10}(0.99\cdot 10^{-3})\simeq 3.004$. We have plotted this number for $N=10^{6}$ in the lower left panel of Figure 2. The simulation-based moments based on $N=1,000,000$ draws from the distribution are far less accurate. At best, the first few decimal places are accurate, and the accuracy deteriorates rapidly as $r$ increases. In the lower right panels we have shown the moments based on the CMGF method for $3.8\leq r\leq 4.2$ represented by the solid blue line. The $\times$-crosses are simulation-based moment estimates, each based on $N=1,000,000$ random draws.
@@
$r$ CMGF Integrate with density Simulations $(N=10^{6})$
@@
4.0 17.6 100.3 318,081
Note: Computation time (in microseconds) for evaluating $\mathbb{E}|X|^{r}$ using three methods: the CMGF method of Theorem 1, by numerical integration of $\int_{-\infty}^{\infty}|x|^{r}f(x)\mathrm{d}x$, and by simulations, where we generate $X_{1},\ldots,X_{N}$ independent and identically NIG distributed, with $N=$1,000,000 and take the average of $|X_{i}|^{r}$. Both CMGF and the density-based method use numerical integration with the same tolerance threshold. Computation times are evaluated with BenchmarkTools.jl for Julia, see Chen and Revels, (2016). Computation were done with Julia v1.11.0, see Bezanson et al., (2017), on a MacBook Pro M1 Max with 32 GM memory.
The CMGF method is not only far more accurate in this example, it is also much faster than the simulation-based approach. This is shown in Table 1 where we report the computation time obtained with a Julia implementation. In this case the density is readily available, and we also compute the moment by evaluating $\int_{-\infty}^{\infty}|x|^{r}f(x)\mathrm{d}x$ using the same numerical integration method used to evaluate the integral in Theorem 1. The CMGF method is between 5 and 25 times faster than the standard approach, and the CMGF method is more than 10,000 times faster than averaging (and generating) $N=10^{6}$ independent draws from the NIG distribution. We have also compared the methods using a standard normal distribution, for which the true moment is known to be, $\mathbb{E}|X|^{r}=\frac{\Gamma(\frac{r+1}{2})}{\sqrt{\pi}}2^{r/2}$, for all $r>-1$. For the Gaussian case the CMGF is also thousands of times faster than simulation methods using 1 million draws and far more accurate. For $|r|<1$, the CMGF is also faster and more accurate than numerical integration based on the density (even if we take advantage of the density being symmetric about zero). For moments larger than one, integration with the density is about twice as fast as CMGF, after taking advantage of the symmetric density.
@@ 3.1 Moments of Cumulative Returns in a GARCH Model
where $r_{f}$ is the risk free rate; $\lambda$ is the equity risk premium; and $h_{t+1}={\rm var}\left(R_{t+1}|\mathcal{F}_{t}\right)$ is the conditional variance of daily log-returns $r_{t+1}=\log\left(S_{t+1}/S_{t}\right)$.
Among the many variations of GARCH models, the HNG model stands out as having an analytical MGF for cumulative returns, $R_{T,H}=\sum_{t=T+1}^{T+H}r_{t}$. The underlying reason is that its dynamic structure is carefully crafted for the purpose of yielding closed-form option pricing formulae, and option pricing formulae depend on the properties of cumulative returns, $R_{T,H}=\sum_{t=T+1}^{T+H}r_{t}$. The conditional density function for $R_{T,H}$ does not have a known analytical form, but the corresponding MGF is available in closed form [^5] We follow standard terminology and label an expression as “closed form” if it can be expressed in terms of elementary mathematical functions and operations, which is the case for our recursive expressions. from simple recursive expressions, as stated below.
@@
Let $r_{t}$, $t=1,2,\ldots$, be given by (7) and define cumulative returns, $R_{T,H}=r_{T+1}+\cdots+r_{t+H}$. Then the conditional MFG has the affine form,
@@
We will illustrate this with a simulation design that is based on real data. We estimate the HNG model using daily log returns for the S&P 500 index from January 1, 2000 to December 30, 2021. The data were obtained from WRDS and the maximum likelihood estimates are presented in Table 2 along with their standard errors (in parentheses).

### v4

@@ 2 A New Method for Computing Moments (CMGF)
Suppose $\mathbb{E}[e^{\pm sX}]<\infty$ for some $s>0$. Then
$\mathbb{E}[(X-\xi)^{k}]$ $=$ $\frac{k!}{\pi}\int_{0}^{\infty}\operatorname{Re}\left[\frac{e^{-\xi z}M_{X}(z)+(-1)^{k}e^{\xi z}M_{X}(-z)}{z^{k+1}}\right]\mathrm{d}t,\quad\text{for }k\in\mathbb{N}_{0}.$ (10)
Because $k$ is real, we have expressed the integral in the form, $\frac{k!}{\pi}\int_{0}^{\infty}\operatorname{Re}[\cdot]\mathrm{d}t$, which tends to be simpler to evaluate numerically than the equivalent expression using the form, $\frac{k!}{2\pi}\int_{-\infty}^{+\infty}[\cdot]\mathrm{d}t$.
@@
In some situations, Theorems 1, 2, and 3, provide different expressions for the same moments. For instance, (4) and (9) must agree for a non-negative random variable, because $\mathbb{E}|X|^{r}=\mathbb{E}[X^{r}]$, and for integer moments, these expressions must also agree with (10). This is indeed the case, because the additional term in (4) and the additional term in (10) are both zero for non-negative variables. This is an implication of Lemma 1, which establishes that $\int_{-\infty}^{\infty}\frac{e^{-\left(s+it\right)x}}{\left(s+it\right)^{r+1}}\mathrm{d}t=0$ for all $x>0$ and all $s>0$.
Negative tail moments, which is part of the definition of expected shortfall (see below), have the following expression, where we can combine Theorems 2 and 3 and use the notation $x_{-}=\min(x,0)$.
Corollary 2 (Negative tail moments).
Suppose $\mathbb{E}[e^{-sX}]<\infty$ for some $s>0$ and $\mathbb{E}|X-\xi|^{k}<\infty$. Then
$\mathbb{E}[(X-\xi)_{-}^{k}]=-\frac{k!}{\pi}\int_{0}^{\infty}{\rm Re}\left[\frac{e^{\xi z}M_{X}(-z)}{z^{k+1}}\right]\mathrm{d}t,\quad\text{for }k\text{ odd}.$
We can also compute the cumulative distribution function (cdf), quantiles, and expected shortfall using the new moment expressions. For a continuous random variable, $X$, its cdf is explicitly given by, $F_{X}(x)=\frac{1}{2}-\frac{1}{\pi}\int_{0}^{\infty}{\rm Re}\left[\frac{M_{X}(it)e^{-itx}}{it}\right]\mathrm{d}t,$ which is the classical Gil-Pelaez inversion formula that provides a direct way to recover probability distributions from their transforms. Here we have used that $M_{X}(it)$ is the characteristic function for $X$. If unique, the $\alpha$-quantile is now given by $\xi_{\alpha}\equiv F_{X}^{-1}(\alpha)$, and we can use Corollary 2 with $k=1$ to compute expected shortfall,
$\operatorname{ES}_{\alpha}(X)=-\frac{1}{\alpha}\mathbb{E}[X1_{\{X<\xi_{\alpha}\}}]=-\frac{1}{\alpha}\mathbb{E}[(X-\xi_{\alpha})1_{\{X<\xi_{\alpha}\}}]-\xi_{\alpha}.$
@@ 2.1 Example: Moments of NIG Distribution
To illustrate the new method, we use Theorem 1 to compute fractional moments of the normal-inverse Gaussian (NIG) distribution. The NIG distribution was introduced in Barndorff-Nielsen, (1978) and has four parameters, $\mu$ (location), $\delta$ (scale), $\alpha$ (tail heaviness), and $\beta$ (asymmetry). It has density
@@
$f(x)=\tfrac{\alpha\delta}{\pi\sqrt{\delta^{2}+(x-\mu)^{2}}}K_{1}\left(\alpha\sqrt{\delta^{2}+(x-\mu)^{2}}\right)e^{\delta\gamma+\beta(x-\mu)},\qquad x\in\mathbb{R}$
where $\gamma=\sqrt{\alpha^{2}-\beta^{2}}$ and $K_{1}(\cdot)$ is the modified Bessel function of the second kind, and the MGF is given by $M_{X}(z)=\exp\left(\mu z+\delta\left[\gamma-\sqrt{\alpha^{2}-(\beta-z)^{2}}\right]\right),$ for $z=s+it$ with $\beta-\alpha<s<\beta+\alpha$.
Existing expressions for fractional moments of NIG distributions are rather complicated. Let $X\sim\mathrm{NIG}(\alpha,\beta,\mu,\delta)$ and $Y\sim\mathrm{NIG}(\alpha,0,\mu,\delta)$, then Barndorff-Nielsen and Stelzer, (2005) showed that
$\mathbb{E}|X|^{r}=e^{\delta(\gamma-\alpha)}\sum_{k=0}^{\infty}\frac{\beta^{k}}{k!}\mathbb{E}\left(|Y|^{r}(Y-\mu)^{k}\right),\qquad r>0,$
which expresses $\mathbb{E}|X|^{r}$ as an infinite sum of moments involving the symmetric NIG, $Y$. For $\mu$-centered moments Barndorff-Nielsen and Stelzer, (2005, corollary 3) obtained the following explicit formula,
$\mathbb{E}(|X-\mu|^{r})=\frac{\alpha}{\pi}\left(\frac{2\delta}{\alpha}\right)^{\tfrac{r+1}{2}}e^{\delta\gamma}\sum_{k=0}^{\infty}\frac{2^{k}\Gamma(k+\tfrac{r+1}{2})}{(2k)!}\left(\frac{\delta\beta^{2}}{\alpha}\right)^{k}K_{k+\tfrac{r-1}{2}}(\delta\alpha),\quad r>0.$ (11)
The corresponding CMGF integral is
$\mathbb{E}|X-\mu|^{r}=\frac{\Gamma(r+1)}{2\pi}e^{\delta\gamma}\int_{-\infty}^{+\infty}\frac{e^{-\delta\sqrt{\alpha^{2}-(\beta-z)^{2}}}+e^{-\delta\sqrt{\alpha^{2}-(\beta+z)^{2}}}}{z^{r+1}}\mathrm{d}t,\qquad r>-1,$
and Theorem 1 yields similar expressions for any other recentering ($\xi\neq\mu$), which is not possible with (11). Evaluating the CMGF integral is also substantially faster and more accurate than both simulation-based methods and direct integration using the NIG density.
@@
To illustrate this, we consider two standardized NIG distribution (zero mean and unit variance), which can be characterized by the two parameters, $(\xi,\chi)$, for $0\leq|\chi|<\xi<1$, where $\mu=-\tfrac{\chi}{\xi}\zeta$, $\delta=\xi^{2}\sqrt{\xi^{2}-\chi^{2}}\zeta$, $\alpha=\xi\zeta$, and $\beta=\chi\zeta$, with $\zeta=\sqrt{1-\xi^{2}}/(\xi^{2}-\chi^{2})$, see Barndorff-Nielsen et al., (1985). Specifically we consider $(\xi,\chi)=(1/2,-1/3)$ and $(\xi,\chi)=(1/8,-1/16)$, the corresponding densities are shown in the left panel of Figure 1. The suitable range for $s$ is $0<s<\alpha+\beta$, where the upper limit for these distributions is $3\sqrt{3}/5\approx 1.04$ and $2\sqrt{63}/3\approx 5.29$, respectively. So, we can set $s=1$ and the mapping, and $t\mapsto\operatorname{MGF}(1+it)$ is shown for the two distributions in the right panel – solid lines for the real part and dashed lines for the imaginary part.
Figure 1: Densities of two standardized NIG distributions (left panel) and the $t\mapstoMGF$($1+it)$ (right panel). The two distributions are for $(\xi,\chi)=(1/2,-1/3)$ (blue lines) and $(\xi,\chi)=(1/8,-1/16)$ (red lines).
@@
We compute the absolute moments for $0.85\leq r\leq 4.2$ for the two distributions using Theorem 1 and the moments are shown in the upper panel of Figure 2. The solid black dot is the second (absolute) moment, which is known to be one for these standardized distributions, and the solid blue and red dots represent the fourth (absolute) moment, which are known to equal $3(1+4\chi^{2})/(1-\xi^{2})$. The values provided by the CMGF method are very accurate. The first 12 digits of the fourth moments are correct using a simple implementation in Julia.
@@
Figure 2: Absolute moments of the standardized NIG distribution with $(\xi,\chi)=(1/2,-1/3)$ (blue lines) and $(\xi,\chi)=(1/8,-1/16)$ (red lines). These are shown in the left panels, where dots represent the known moments for $r=2$ and $r=4$. The middle panel is a snippet of the left panel where we have added x-crosses that represents 100 simulation-based estimates of $\mathbb{E}|X|^{r}$, for $r\in\{3.82,3.84,\ldots,4.18\}$. Each estimate is based on N=N=1,000,000 random draws of the NIG distribution. The right panel shows how many decimal places are accurate with a one standard deviation simulation error.
@@
Additionally, we estimate the absolute moments using $N=1,000,000$ independent draws from the two NIG distributions. The simulation-based moment is given by $\hat{\mu}_{N}^{r}=\frac{1}{N}\sum_{j=1}^{N}|X_{i}|^{r}$ and we estimate the $\sigma^{2}(r)=\operatorname{var}(|X_{j}|^{r})$ with 100 million draws. The accuracy of a simulated moment is quantified by its standard error. Assuming an error equal to one standard deviation, then the number of accurate decimal places is given by $\log_{10}(\sigma(r)/\sqrt{N})$. [^8] For instance, a standard error equal to $0.00099$ and $N=10^{6}$ will map to $\log_{10}(0.99\cdot 10^{-3})\simeq 3.004$. We have plotted this number for $N=10^{6}$ in the lower left panel of Figure 2. The simulation-based moments based on $N=1,000,000$ draws from the distribution are far less accurate. At best, the first few decimal places are accurate, and the accuracy deteriorates rapidly as $r$ increases. In the lower right panels we have shown the moments based on the CMGF method for $3.8\leq r\leq 4.2$ represented by the solid blue line. Each $\times$-cross is a simulation-based moment estimate based on $N=1,000,000$ random draws.
@@
$r$ CMGF $M_{X}^{(k)}$ Integrate with density Simulations $(N=10^{6})$
@@
4.0 17.6 104.9 100.3 318,081
Note: Computation time in microseconds (µs) for evaluating $\mathbb{E}|X|^{r}$ using four methods: the CMGF method of Theorem 1, the $k$-th derivative of the MGF (for $r=2$ and $r=4$), by numerical integration of $\int_{-\infty}^{\infty}|x|^{r}f(x)\mathrm{d}x$, and by simulations, where we generate $X_{1},\ldots,X_{N}$ independent and identically NIG distributed, with N=N=1,000,000 and take the average of $|X_{i}|^{r}$. Both CMGF and the density-based method use numerical integration with the same tolerance threshold. Computation times are evaluated with BenchmarkTools.jl for Julia, see Chen and Revels, (2016). Computations were done with Julia v1.11.0, see Bezanson et al., (2017), on a MacBook Pro M1 Max with 32 GM memory.
The CMGF method is not only far more accurate in this example, it is also much faster than the simulation-based approach. This is shown in Table 1 where we report the computation time obtained with a Julia implementation. In this case the density is readily available, and we also compute the moment by evaluating $\int_{-\infty}^{\infty}|x|^{r}f(x)\mathrm{d}x$ using the same numerical integration method used to evaluate the integral in Theorem 1. The CMGF method is between 5 and 25 times faster than the standard approach, and the CMGF method is more than 10,000 times faster than averaging (and generating) $N=10^{6}$ independent draws from the NIG distribution. We have also compared the methods using a standard normal distribution, for which the true moment is known to be, $\mathbb{E}|X|^{r}=\frac{\Gamma(\frac{r+1}{2})}{\sqrt{\pi}}2^{r/2}$, for all r>1r>-1. For the Gaussian case the CMGF is also thousands of times faster than simulation methods using 1 million draws and far more accurate. For $|r|<1$, the CMGF is also faster and more accurate than numerical integration based on the density (even if we take advantage of the density being symmetric about zero). For moments larger than one, integration with the density is about twice as fast as CMGF, after taking advantage of the symmetric density.
@@ 3.1 Moments of Cumulative Returns in a GARCH Model
where $r_{f}$ is the risk-free rate; $\lambda$ is the equity risk premium; and $h_{t+1}={\rm var}\left(R_{t+1}|\mathcal{F}_{t}\right)$ is the conditional variance of daily log-returns $r_{t+1}=\log\left(S_{t+1}/S_{t}\right)$.
Among the many variations of GARCH models, the HNG model stands out as having an analytical MGF for cumulative returns, $R_{T,H}=\sum_{t=T+1}^{T+H}r_{t}$, where $H$ is the horizon in trading days. The underlying reason is that its dynamic structure is carefully crafted for the purpose of yielding closed-form option pricing formulae, and option pricing formulae depend on the properties of cumulative returns, $R_{T,H}=\sum_{t=T+1}^{T+H}r_{t}$. The conditional density function for $R_{T,H}$ does not have a known analytical form, but the corresponding MGF is available in closed form [^9] We follow standard terminology and label an expression as “closed form” if it can be expressed in terms of elementary mathematical functions and operations, which is the case for our recursive expressions. from simple recursive expressions, as stated below.
@@
Let $r_{t}$, $t=1,2,\ldots$, be given by (12) and define cumulative returns, $R_{T,H}=r_{T+1}+\cdots+r_{t+H}$. Then the conditional MGF has the affine form,
@@
We will illustrate this with a simulation design that is based on real data. We estimate the HNG model using daily log-returns for the S&P 500 index from January 1, 2000 to December 30, 2021. The data were obtained from WRDS and the maximum likelihood estimates are presented in Table 2 along with their standard errors (in parentheses).